for a convex lens 1/v+1/u=1/f .  how u would determine focal length graphicallyfor a convex lens 1/v+1/u=1/f .  how u would determine focal length graphicallyU

Grade:

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Rishabh,

Ans:- As for a particular lens, the focal length is fixed hence f is fixed and hence 1/f is also fixed.

So, lets take 1/v=x and 1/u=y So, in the experiment take various object and image distances and plot their inverse in the graph paper.

So we shall get, a st line x+y=k where K =1/f. Now transfer it is the form of x/k + y/k=1 from the plotted st line and it,s intercept we can find out the value of k and then find out

it's inverse which will give the value of the focal length 

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Rishabh!!!

 


Regards,

Askiitians Experts

SOUMYAJIT IIT_KHARAGPUR